# Towards Joint Optimization for DNN Architecture and Configuration for Compute-In-Memory Hardware

### 19 Feb 2024 v1

@@ Abstract
1 Introduction
The Von-Neumann architecture von Neumann (1945) is widely adopted with several advancements Johnson and Johnson (1991); Gochman et al. (2006) across applications and domains. However, for data-intensive requirements as prevalent in DNN models Brown et al. (2020); Touvron et al. (2023), it suffers from memory bandwidth and on-chip interconnect limitations. On the other hand, recent technological advancements such as 3D and 2.5D stacking, blending of logic and memory in the process have enabled Compute-in-Memory (CiM) to provide unprecedented benefits Singh et al. (2019); Fayyazi et al. (2019).
@@ 1 Introduction
Figure 1: (a) Illustration of the proposed joint architecture-hardware configuration search framework. (b) Illustration of different possible positions for the placement of a dedicated and specialized compute unit to perform MAC operations.
Neural architecture search (NAS) Zoph and Le (2016); Zoph et al. (2018); Cai et al. (2019) has emerged as a popular method for selection of near-optimal sub-network architectures that can be run on a given hardware platform efficiently while attaining a high task accuracy. NAS can yield a range of Pareto-optimal sub-networks that can provide efficient trade-offs between accuracy and efficiency.
Based on the training frequency of a network on various use-cases across diverse platforms, NAS can be divided into two broad categories, namely iterative Cai et al. (2018) and one-shot NAS Cai et al. (2019); Sridhar et al. (2023); Zhang et al. (2023). As the name suggests, in the iterative approach, the network needs to be retrained for any change in the underlying hardware constraints. One-shot NAS on the other hand, trains a super-network only once, and samples suitable sub-networks based on different hardware constraints associated to different devices.
@@
Our Contribution. With the slow-down of Moore’s law, the right choice of hardware configuration to yield efficient DNN architectures has become important in improving the model’s hardware efficiency. Towards that goal, we present CiMNet, a joint optimization framework in yielding not only sub-network architectures but also hardware configurations to get superior model-configuration choices. Specifically, CiMNet performs a joint search over the model elastic parameters as well as the CiM hardware elastic configurations [^1] In this paper, we refer to elastic hardware variables as ”configurations” and elastic model parameters as ”architecture” choices. to evaluate the accuracy and cycles for various sub-networks. This is performed using a multi-objective evolutionary search to yield choices of various sub-networks and associated hardware configurations that can improve accuracy while requiring significantly fewer cycles. Unlike the few existing approaches of joint hardware-model architecture search Hao et al. (2019) that rely on a sequential and time consuming iterative approach, we present a parallel selection of both the architecture and configuration choices. Additionally, as depicted in Figure 1(a), we leverage a predictor-based approach to estimate the accuracy and cycles of a large corpus of sub-networks and associated hardware configurations. We thus avoid real evaluations for each sub-network, allowing for a significant speed up of the search process. Additionally, to achieve design space exploration and its implication on model performance, we develop an infrastructure for a DNN graph compiler and a cycle-accurate simulator for CiM hardware which can project end-to-end latency and data-transfers for any given hardware configuration and sub-network.
@@
2 Background and Motivation
2.1 CiM Accelerators for DNNs
In CiM architectures the MAC operation is performed in the memory array either by leveraging bitline current or voltage (bitline compute) or placing a specialised compute unit in the peripheral circuitry (bitline-free compute) Ramanathan et al. (2020a).
In the bitline compute approach, MAC operations are performed inside the memory array by leveraging Kirchhoff’s circuit laws (marked as 1 in Figure 1 (b)). Typically, these architectures leverage crossbar memory arrays to store weights and perform the MAC operations by exciting multiple rows (word-lines of the crossbar array) and produce the output by performing a read operation on each column (bit-lines of the crossbar array) Chi et al. (2016); Shafiee et al. (2016); Joshi et al. (2020); Le Gallo et al. (2023).
In the bitline-free compute approach Ramanathan et al. (2020b), peripheral circuits are leveraged to place a dedicated specialized compute unit for each memory array to perform the MAC operation (marked as 2 in Figure 1(b)) Ramanathan et al. (2020a). Unlike in bitline compute approach, conventional memory read and write operations are used to access and store the data. This approach requires no modifications to the memory array but some modification is required to accommodate a specialized compute unit in the periphery.
Other works Chen et al. (2017); Jouppi and et al. (2017) that integrate specialized processing unit for MAC operations and a memory unit on the same silicon (marked as 3 in Figure 1(b)) are not classified CiM. Our experiments and evaluations of the joint-optimization approach are performed on a bitline-free CiM architecture. Nonetheless, our approach should scale to other types of hardware architectures such as Von-Neumann and FPGA as well.
2.2 CiM Architecture
@@
Figure 2: Illustration of the compute in-memory architecture used in our experiments and evaluation of joint-search.
3 Related Work
3.1 Hardware Design Space Optimization
@@
Researchers have explored several different methods for optimizing hardware resources and usage to improve execution performance. In Mehrabi et al. (2021), the authors explored Bayesian optimization for improving execution latency and resource usage on FPGAs. In Ascia et al. (2007), the authors have demonstrated the use of multi-objective evolutionary algorithms for designing optimal application specific system-on-chip and fuzzy algorithms for estimating hardware performance metrics for multimedia applications. Researchers in Ascia et al. (2006) have explored genetic algorithms for the efficient design of embedded systems. Similarly, in Eyerman et al. (2006) authors have explored the design of out-of-order embedded processors using genetic algorithms. In Dubach et al. (2007), the authors explored predictor-guided search techniques for the micro-architectural design space exploration to optimize the energy-delay product. To the best of our knowledge, this work is the first attempt at design exploration of CiM hardware.
3.2 Neural Architecture Search
@@
NAS is a popular method of optimized model architecture search for a given hardware constraint Zoph et al. (2018); Cai et al. (2018); Kundu et al. (2021, 2023, 2022). Initially, NAS methods Zoph and Le (2016) focused on accuracy rather than their hardware deployment efficiency including FLOPs and cycles. Later hardware-aware NAS methods Cai et al. (2019); Wu et al. (2019); Stamoulis et al. (2019) optimized model search by directly providing the hardware feedback into architecture search. Hardware-DNN co-design techniques Hao et al. (2019); Jiang et al. (2019) jointly optimize neural network architectures and hardware configurations. Specifically, this work performs hardware-oriented DNN model search, then leverages an Auto-HLS engine to generate synthesizable C code of the FPGA accelerator for explored DNNs. As a result, they can improve inference efficiency. However, given new inference hardware platforms, these methods need to repeat the architecture search process as well as perform iterative training. This essentially leads to prohibitively high computational time as well as carbon emissions. Moreover, the sequential nature of the joint exploration (search over architecture first and then leverage tools like AutoHLS Hao et al. (2019) to yield optimal hardware simulation) make the hardware design space choice a function of a small corpus of sub-networks. Moreover, such approaches are not scalable to a large number of deployment scenarios. Recently another line of work Lin et al. (2020) has leveraged one-shot NAS to perform joint architecture and efficient runtime search on micro controllers. This work performed joint compiler hyperparameter (e.g., TVM) and architecture co-search to better tune a compiler for a given architecture. Despite being efficient, such approaches have yet to handle the changing hardware configuration parameters. Moreover, the majority of the one-shot NAS work assumes to have a look-up table of pre-computed hardware metrics for an underlying configuration, which is not possible to have for a dynamically changing hardware configuration.
4 Methods
@@
4.1 Hardware Configuration Parameters
@@
4.2 CiMNet: Joint Optimization Framework
To perform search with both neural network architectures and hardware configurations, we start an iteration by randomly sampling a set of sampled elastic architecture and configuration choices. We then represent each selected elastic model architecture and hardware configuration choice in a one-hot encoded way. In particular, for hardware configurations, we encode eight elastic variables as shown in Table 1. For DNN architecture we encode kernel size, channel width factor, and depth factor as elastic variable for convolutional architectures. For ViT, we use the number of layers, number of attention heads and intermediate dimension as the elastic architecture choices. Once the one-hot encoded joint configuration inputs are created, we use them to create the input vectors that are used to train a accuracy and a cycles predictor. For the accuracy predictor we compute the true accuracy for a small set of sub-networks by running their validation performance. Similarly, for the cycle predictor, we use the chosen hardware configurations and sub-networks to evaluate the true cycles via simulation.
We then use trained predictors to predict cycles and accuracy for a large corpus of selected samples $M$ that are necessary to evaluate during the next step of our method, a multi-objective (accuracy and cycle count, here) search. We use NSGA-II Yusoff et al. (2011) evolutionary search for this step. Note the use of predictors dramatically reduces the search time. After the search is complete, we select the top-$N$ networks to determine their true accuracy and cycle count. Here, $M>>N$, and thus the predictor based approach for the multi-objective search speeds up ranking operation by around $\frac{M}{N}$ times as prediction takes negligible time to estimate values. This also makes our approach scalable for search over a large corpus of samples which can otherwise be extremely time-consuming in alternate approaches of hardware evaluations that are present in literature. The training sets of accuracy and cycle count predictors are then updated using these latest $N$ inputs at the end of the current iteration. Thus, at the end of each iteration, the data set used to train the predictors keep on increasing by $N$ thereby improving predictor performance in the following iteration. So, essentially for $K$ iterations the predictor gets a training set of $KN$ samples to train each predictor.
@@ 4.2 CiMNet: Joint Optimization Framework
Table 1: Hardware configurations that are considered elastic for our joint search framework.
4.3 Predictors
Since evaluations of performance objectives, such as top-1 accuracy, require a large amount of time, we take inspiration from Cai et al. (2019) and Wang et al. (2020) and employ predictors. Specifically, we predict the top-1 accuracy and cycle count of the sub-networks sampled from super-networks for a chosen hardware configuration.
Following the same path as Cummings et al. (2022b), we use simpler methods such as ridge regression and support vector machine regression (SVR) predictors, instead of traditional multi-layer perceptron (MLP) based alternative. These simpler methods converge more quickly and require both fewer training examples and fewer hyper-parameter optimization than MLPs Cummings et al. (2022b). Regression is then performed with the one-hot-encoded input vector on the combined search space of accuracy and cycles.
The use of performance objective prediction via simple predictors allows us to significantly accelerate the selection of sub-networks with minimal prediction error Cummings et al. (2022a). Please refer to Section 5.2 for an analysis of cycle count prediction performance.
5 Experimental Results
5.1 Experimental Setup
@@
5.2 Predictor Analysis
As described in Section 4.3, our work makes extensive use of predictors to accelerate the selection of sub-networks. The process used to evaluate the performance of cycle count prediction is the same as described in Cummings et al. (2022a). The averaged results are used to compute both the mean absolute percentage error (MAPE) and correlations.
@@ 5.2 Predictor Analysis
Figure 3: MAPE of predictors performing cycle count prediction versus the number of training examples for sub-networks derived from the MobileNetV3, ResNet-50 and ViT-B super-networks (top row). Correlation and Kendall $\tau$ coefficient between actual and predicted values after training the predictor with 1000 examples (bottom row). Green lines show The ideal correlation.
@@
Table 2: Kendall rank correlation coefficients of different search spaces for sub-networks derived from the MobileNetV3, ResNet-50, and ViT-B super-networks.
5.3 Pareto Front Analysis
@@
Figure 4: Pareto fronts in the accuracy / cycle count space of sub-networks optimized from MobileNetV3, ResNet-50 and ViT-B super-networks for both static (blue) and elastic (green) hardware configurations. The performance when using the static, canonical network architecture along with the static hardware configuration is shown by the black square.
@@
Table 3: Reduction in cycle count for sub-networks derived from the MobileNetV3, ResNet-50 and ViT-B super-networks. These reductions are computed between the baseline cycle count and the optimized cycle count at the baseline accuracy. See Figure 4 for optimized and baseline performance.
@@
Table 4: Hardware configurations parameters and associated normalized values found after joint optimization with MobileNetV3 architectures. $C_{s}$ denotes the static hardware configuration while $C_{min}$, $C_{med}$ and $C_{max}$ denote the elastic hardware configurations associated with the minimum, median and maximum cycle counts respectively. Note that $C_{min}$ and $C_{max}$ correspond to the leftmost and rightmost points respectively of the green Pareto front shown in the first plot of Figure 4.
6 Conclusions
@@ Appendix A Mapping strategy for a deep neural network
Table 5: Dimensions of a DNN layer that are chunked to estimate temporal and spatial dataflows.

### 18 Mar 2024 v2

# CiMNet: Towards Joint Optimization for DNN Architecture and Configuration for Compute-In-Memory Hardware

@@ Abstract.
Compute in memory, NAS, neural architecture and hardware configuration co-search, CNN, Transformer
copyright: rightsretainedjournalyear: 2024
1. Introduction
The Von-Neumann architecture (von Neumann, 1945) is widely adopted with several advancements (Johnson and Johnson, 1991; Gochman et al., 2006) across applications and domains. However, for data-intensive requirements as prevalent in DNN models (Brown et al., 2020; Touvron et al., 2023), it suffers from memory bandwidth and on-chip interconnect limitations. On the other hand, recent technological advancements such as 3D and 2.5D stacking, blending of logic and memory in the process have enabled Compute-in-Memory (CiM) to provide unprecedented benefits (Singh et al., 2019; Fayyazi et al., 2019).
@@ 1. Introduction
Figure 1. (a) Illustration of the proposed joint architecture-hardware configuration search framework. (b) Illustration of different possible positions for the placement of a dedicated and specialized compute unit to perform MAC operations.
Neural architecture search (NAS) (Zoph and Le, 2016; Zoph et al., 2018; Cai et al., 2019) has emerged as a popular method for selection of near-optimal sub-network architectures that can be run on a given hardware platform efficiently while attaining a high task accuracy. NAS can yield a range of Pareto-optimal sub-networks that can provide efficient trade-offs between accuracy and efficiency.
Based on the training frequency of a network on various use-cases across diverse platforms, NAS can be divided into two broad categories, namely iterative (Cai et al., 2018) and one-shot NAS (Cai et al., 2019; Sridhar et al., 2023; Zhang et al., 2023). As the name suggests, in the iterative approach, the network needs to be retrained for any change in the underlying hardware constraints. One-shot NAS on the other hand, trains a super-network only once, and samples suitable sub-networks based on different hardware constraints associated to different devices.
@@
Our Contribution. With the slow-down of Moore’s law, the right choice of hardware configuration to yield efficient DNN architectures has become important in improving the model’s hardware efficiency. Towards that goal, we present CiMNet, a joint optimization framework in yielding not only sub-network architectures but also hardware configurations to get superior model-configuration choices. Specifically, CiMNet performs a joint search over the model elastic parameters as well as the CiM hardware elastic configurations [^1] In this paper, we refer to elastic hardware variables as ”configurations” and elastic model parameters as ”architecture” choices. to evaluate the accuracy and cycles for various sub-networks. This is performed using a multi-objective evolutionary search to yield choices of various sub-networks and associated hardware configurations that can improve accuracy while requiring significantly fewer cycles. Unlike the few existing approaches of joint hardware-model architecture search (Hao et al., 2019) that rely on a sequential and time consuming iterative approach, we present a parallel selection of both the architecture and configuration choices. Additionally, as depicted in Figure 1(a), we leverage a predictor-based approach to estimate the accuracy and cycles of a large corpus of sub-networks and associated hardware configurations. We thus avoid real evaluations for each sub-network, allowing for a significant speed up of the search process. Additionally, to achieve design space exploration and its implication on model performance, we develop an infrastructure for a DNN graph compiler and a cycle-accurate simulator for CiM hardware which can project end-to-end latency and data-transfers for any given hardware configuration and sub-network.
@@
2. Background and Motivation
2.1. CiM Accelerators for DNNs
In CiM architectures the MAC operation is performed in the memory array either by leveraging bitline current or voltage (bitline compute) or placing a specialised compute unit in the peripheral circuitry (bitline-free compute) (Ramanathan et al., 2020a).
In the bitline compute approach, MAC operations are performed inside the memory array by leveraging Kirchhoff’s circuit laws (marked as 1 in Figure 1 (b)). Typically, these architectures leverage crossbar memory arrays to store weights and perform the MAC operations by exciting multiple rows (word-lines of the crossbar array) and produce the output by performing a read operation on each column (bit-lines of the crossbar array) (Chi et al., 2016; Shafiee et al., 2016; Joshi et al., 2020; Le Gallo et al., 2023).
In the bitline-free compute approach (Ramanathan et al., 2020b), peripheral circuits are leveraged to place a dedicated specialized compute unit for each memory array to perform the MAC operation (marked as 2 in Figure 1(b)) (Ramanathan et al., 2020a). Unlike in bitline compute approach, conventional memory read and write operations are used to access and store the data. This approach requires no modifications to the memory array but some modification is required to accommodate a specialized compute unit in the periphery.
Other works (Chen et al., 2017; Jouppi and et al., 2017) that integrate specialized processing unit for MAC operations and a memory unit on the same silicon (marked as 3 in Figure 1(b)) are not classified CiM. Our experiments and evaluations of the joint-optimization approach are performed on a bitline-free CiM architecture. Nonetheless, our approach should scale to other types of hardware architectures such as Von-Neumann and FPGA as well.
2.2. CiM Architecture
@@
Figure 2. Illustration of the compute in-memory architecture used in our experiments and evaluation of joint-search.
3. Related Work
3.1. Hardware Design Space Optimization
@@
Researchers have explored several different methods for optimizing hardware resources and usage to improve execution performance. In (Mehrabi et al., 2021), the authors explored Bayesian optimization for improving execution latency and resource usage on FPGAs. In (Ascia et al., 2007), the authors have demonstrated the use of multi-objective evolutionary algorithms for designing optimal application specific system-on-chip and fuzzy algorithms for estimating hardware performance metrics for multimedia applications. Researchers in (Ascia et al., 2006) have explored genetic algorithms for the efficient design of embedded systems. Similarly, in (Eyerman et al., 2006) authors have explored the design of out-of-order embedded processors using genetic algorithms. In (Dubach et al., 2007), the authors explored predictor-guided search techniques for the micro-architectural design space exploration to optimize the energy-delay product. To the best of our knowledge, this work is the first attempt at design exploration of CiM hardware.
3.2. Neural Architecture Search
@@
NAS is a popular method of optimized model architecture search for a given hardware constraint (Zoph et al., 2018; Cai et al., 2018; Kundu et al., 2021, 2023, 2022). Initially, NAS methods (Zoph and Le, 2016) focused on accuracy rather than their hardware deployment efficiency including FLOPs and cycles. Later hardware-aware NAS methods (Cai et al., 2019; Wu et al., 2019; Stamoulis et al., 2019) optimized model search by directly providing the hardware feedback into architecture search. Hardware-DNN co-design techniques (Hao et al., 2019; Jiang et al., 2019) jointly optimize neural network architectures and hardware configurations. Specifically, this work performs hardware-oriented DNN model search, then leverages an Auto-HLS engine to generate synthesizable C code of the FPGA accelerator for explored DNNs. As a result, they can improve inference efficiency. However, given new inference hardware platforms, these methods need to repeat the architecture search process as well as perform iterative training. This essentially leads to prohibitively high computational time as well as carbon emissions. Moreover, the sequential nature of the joint exploration (search over architecture first and then leverage tools like AutoHLS (Hao et al., 2019) to yield optimal hardware simulation) make the hardware design space choice a function of a small corpus of sub-networks. Moreover, such approaches are not scalable to a large number of deployment scenarios. Recently another line of work (Lin et al., 2020) has leveraged one-shot NAS to perform joint architecture and efficient runtime search on micro controllers. This work performed joint compiler hyperparameter (e.g., TVM) and architecture co-search to better tune a compiler for a given architecture. Despite being efficient, such approaches have yet to handle the changing hardware configuration parameters. Moreover, the majority of the one-shot NAS work assumes to have a look-up table of pre-computed hardware metrics for an underlying configuration, which is not possible to have for a dynamically changing hardware configuration.
4. Methods
@@
4.1. Hardware Configuration Parameters
@@
4.2. CiMNet: Joint Optimization Framework
To perform search with both neural network architectures and hardware configurations, we start an iteration by randomly sampling a set of sampled elastic architecture and configuration choices. We then represent each selected elastic model architecture and hardware configuration choice in a one-hot encoded way. In particular, for hardware configurations, we encode eight elastic variables as shown in Table 1. For DNN architecture we encode kernel size, channel width factor, and depth factor as elastic variable for convolutional architectures. For ViT, we use the number of layers, number of attention heads and intermediate dimension as the elastic architecture choices. Once the one-hot encoded joint configuration inputs are created, we use them to create the input vectors that are used to train an accuracy and a cycles predictor. For the accuracy predictor we compute the true accuracy for a small set of sub-networks by running their validation performance. Similarly, for the cycle predictor, we use the chosen hardware configurations and sub-networks to evaluate the true cycles via simulation.
We then use trained predictors to predict cycles and accuracy for a large corpus of selected samples $M$ that are necessary to evaluate during the next step of our method, a multi-objective (accuracy and cycle count, here) search. We use NSGA-II (Yusoff et al., 2011) evolutionary search for this step. Note, the use of predictors dramatically reduces the search time. After the search is complete, we select the top-$N$ networks to determine their true accuracy and cycle count. Here, $M>>N$, and thus the predictor based approach for the multi-objective search speeds up ranking operation by around $\frac{M}{N}$ times as prediction takes negligible time to estimate values. This also makes our approach scalable for search over a large corpus of samples which can otherwise be extremely time-consuming in alternate approaches of hardware evaluations that are present in literature. The training sets of accuracy and cycle count predictors are then updated using these latest $N$ inputs at the end of the current iteration. Thus, at the end of each iteration, the data set used to train the predictors keep on increasing by $N$ thereby improving predictor performance in the following iteration. So, essentially for $K$ iterations the predictor gets a training set of $KN$ samples to train each predictor.
@@ 4.2. CiMNet: Joint Optimization Framework
Table 1. Hardware configurations that are considered elastic for our joint search framework.
4.3. Predictors
Since evaluations of performance objectives, such as top-1 accuracy, require a large amount of time, we take inspiration from (Cai et al., 2019) and (Wang et al., 2020) and employ predictors. Specifically, we predict the top-1 accuracy and cycle count of the sub-networks sampled from super-networks for a chosen hardware configuration.
Following the same path as (Cummings et al., 2022b), we use simpler methods such as ridge regression and support vector machine regression (SVR) predictors, instead of traditional multi-layer perceptron (MLP) based alternative. These simpler methods converge faster and require both fewer training examples and fewer hyper-parameter optimization than MLPs (Cummings et al., 2022b). Regression is then performed with the one-hot-encoded input vector on the combined search space of accuracy and cycles.
The use of performance objective prediction via simple predictors allows us to significantly accelerate the selection of sub-networks with minimal prediction error (Cummings et al., 2022a). Please refer to Section 5.2 for an analysis of cycle count prediction performance.
5. Experimental Results
5.1. Experimental Setup
@@
5.2. Predictor Analysis
As described in Section 4.3, our work makes extensive use of predictors to accelerate the selection of sub-networks. The process used to evaluate the performance of cycle count prediction is the same as described in (Cummings et al., 2022a). The averaged results are used to compute both the mean absolute percentage error (MAPE) and correlations.
@@ 5.2. Predictor Analysis
Figure 3. MAPE of predictors performing cycle count prediction versus the number of training examples for sub-networks derived from the MobileNetV3, ResNet-50 and ViT-B super-networks (top row). Correlation and Kendall $\tau$ coefficient between actual and predicted values after training the predictor with 1000 examples (bottom row). Green lines show The ideal correlation.
@@
Table 2. Kendall rank correlation coefficients of different search spaces for sub-networks derived from the MobileNetV3, ResNet-50, and ViT-B super-networks.
5.3. Pareto Front Analysis
@@
Figure 4. Pareto fronts in the accuracy / cycle count space of sub-networks optimized from MobileNetV3, ResNet-50 and ViT-B super-networks for both static (blue) and elastic (green) hardware configurations. The performance when using the static, canonical network architecture along with the static hardware configuration is shown by the black square.
@@
Table 3. Reduction in cycle count for sub-networks derived from the MobileNetV3, ResNet-50 and ViT-B super-networks. These reductions are computed between the baseline cycle count and the optimized cycle count at the baseline accuracy. See Figure 4 for optimized and baseline performance.
@@
Table 4. Hardware configurations parameters and associated normalized values found after joint optimization with MobileNetV3 architectures. $C_{s}$ denotes the static hardware configuration while $C_{min}$, $C_{med}$ and $C_{max}$ denote the elastic hardware configurations associated with the minimum, median and maximum cycle counts respectively. Note that $C_{min}$ and $C_{max}$ correspond to the leftmost and rightmost points respectively of the green Pareto front shown in the first plot of Figure 4.
6. Conclusions
@@ Appendix A Mapping strategy for a deep neural network
Table 5. Dimensions of a DNN layer that are chunked to estimate temporal and spatial dataflows.